# Required-edge Cycle Cover Problem: an ASP-Completeness Framework for Graph Problems and Puzzles

### 03 Mar 2026 v1

@@ 1  Introduction
Another framework is the Constraint Graph Satisfiability problem (CGS) introduced in [HearnDemaine09Puzzles, Section 5.1.3], which is a variant of the graph orientation problem defined on constraint graphs. In essence, a constraint graph is a graph with edge weights $1$ or $2$, where the vertices are of three types: and, or, and maj (majority). While planar CGS with only and and or vertices is known to be NP-complete [HearnDemaine09Puzzles, Theorem 5.4], extending this result to ASP-completeness is nontrivial. This is because the known reduction relies on a red-blue conversion gadget ([HearnDemaine09Puzzles, Figure 2.4]) that is not parsimonious, and constructing a parsimonious version is conjectured to be impossible [MITHardness24GeneralizedSAT]. To address this distinction, recent work [MITHardness24GeneralizedSAT] formalized two variants of CGS: arbitrary edge weights, where an edge may have different weights at its two endpoints, and the standard matching edge weights, where an edge must have the same weight at both endpoints; the former is essentially equivalent to a red-blue conversion. Regarding the restriction of vertex types, the complexity classification for CGS with arbitrary edge weights is now fully settled [MITHardness24GeneralizedSAT]; in particular, the case with and, or, and maj vertices is known to be ASP-complete even when the graph is planar. In contrast, the ASP-completeness of CGS with matching edge weights using and, or, and maj vertices has remained open.
@@ 1.1  Our Contributions
In this paper, we propose an ASP-completeness framework designed to exploit the rich structural properties of puzzles, and demonstrate its versatility by applying it to several puzzle types. Specifically, our framework is based on a variant of the Cycle Cover Problem (CCP), which asks whether a given graph can be covered by vertex-disjoint cycles. CCP on mixed graphs (i.e., graphs with both directed and undirected edges) was originally introduced to analyze the ASP-completeness of Kakuro (Cross Sum) puzzles, and was shown to be ASP-complete even when restricted to planar max-degree-$3$ graphs [Seta02CrossSum]. First, we introduce required edges—edges that must be included in the cycle cover—into CCP on mixed graphs, and define this problem as the Required-edge Cycle Cover Problem (RCCP). We focus on a subclass of instances on planar degree-$3$ bipartite graphs where each vertex is incident to exactly one required edge; we refer to this problem as Restricted RCCP. We prove that Restricted RCCP is ASP-complete. Moreover, by constructing a parsimonious gadget simulating a required edge, we obtain a corollary for CCP restricted to planar degree-$3$ bipartite mixed graphs where each vertex is incident to at least one directed edge. We denote this variant as Restricted CCP and conclude that it is also ASP-complete.
Second, we establish the ASP-completeness of several problems via reductions from Restricted RCCP/CCP. Crucial to these reductions is the exclusive structure of Restricted RCCP/CCP: since every vertex is of degree $3$, fixing a cycle cover enforces a strict local constraint where each vertex must have exactly one incoming edge, one outgoing edge, and one unused edge. From the edge perspective, each edge can take one of three states: being traversed in either direction or remaining unused (with directed edges limited to two states). Since each consistent assignment of these states corresponds to a unique cycle cover, this structure allows us to simulate Restricted RCCP/CCP using simple vertex and edge gadgets. Furthermore, the bipartiteness of the graphs permits the use of asymmetric edge gadgets and facilitates the geometric embedding of these gadgets into the rectangular grids typical of puzzles.
@@ 2  Preliminaries
A mixed graph $G=(V,E,A)$ is a graph consisting of a set of vertices $V$, a set of undirected edges $E$, and a set of directed edges $A$. The degree of a vertex $v$ is defined as the number of undirected edges incident to $v$ plus the number of directed edges incident to $v$ (i.e., having $v$ as a head or tail).
@@ 3.2  Flow Model Equivalent to Restricted RCCP
In order to construct the model, we assume that each source has an in-degree of 0, an out-degree of $2$, and a supply of $2$. Under this assumption, an unconstrained source admits any flow distributions in $\{(0,2),(1,1),(2,0)\}$. We additionally employ three variants that restrict the admissible flow distributions as follows:
@@
$\{(1,1),(2,0)\}$. Must send at least $1$ unit along the first edge. We say such a source is oriented toward that edge.
@@
Formally, we transform an instance of Restricted RCCP into a flow network as follows. First, given a planar degree-$3$ mixed graph $G$ with required edges, we rectilinearly embed it into a grid (Figure˜4(a)). Let $H$ denote the resulting grid graph containing the embedding (Figure˜4(b)). We construct a network $N$ from $H$ based on the following rules (Figure˜4(c)):
@@
    Each segment (including empty ones) of $H$ is subdivided by inserting a source of supply $2$; the two resulting edges are oriented toward the adjacent sinks.
@@
Now, we associate cycle covers of $G$ with flows in $N$ as follows. For each vertex $v$ and its corresponding sink, an edge incident to $v$ is incoming, outgoing, or unused in the cycle cover if and only if the corresponding flow value is $2$, 0, or $1$, respectively. Under this correspondence, we have the following lemma.
@@
lemmalemflowmodel Let $G$ be a planar degree-$3$ mixed graph in which every vertex is incident to exactly one required edge, and let $N$ be the flow network constructed as described above. Then there is a bijection between the cycle covers of $G$ and the flows in $N$.
Proof.We first verify the forward direction. Fix a cycle cover of $G$, and consider its corresponding cycle family in $H$. Each lattice point in $H$ is either traversed by the cycle family or not.. In the former case, the sources corresponding to the incoming and outgoing edges of the cycle contribute $2$ and 0 units of flow, respectively; each source on the remaining incident edges contributes $1$ unit. In the latter case, all incident edges are unused; thus each source contributes $1$ unit of flow. Therefore, the total inflow to the sink equals its degree in $H$ in any case.
Conversely, we show that a feasible flow in $N$ induces a valid cycle cover of $G$. Fix a feasible flow in $N$. We construct a subgraph of $G$ and orient its edges based on the flow interpretation defined above. Flow conservation at intermediate lattice points ensures that each edge of $G$ has a well-defined orientation. Furthermore, due to the biased sources, this orientation is consistent with the direction of the directed edges in $G$.
The remaining task is to show that this subgraph forms a valid cycle cover. Consider a vertex $v$ in $G$ and its corresponding sink $s$ in $N$. Assume first that $s$ has degree $3$. Since $v$ has degree $3$ and is incident to exactly one required edge, one of the three sources adjacent to $s$ is an exclusive source, which supplies either 0 or $2$ units of flow. Since the flow from each source is at most $2$, the only way that three nonnegative integers, one of which is in $\{0,2\}$, can sum to $3$ is for the multiset of contributions to be $\{0,1,2\}$. The case where $s$ has degree $4$ is similar. Indeed, exactly three of its adjacent sources correspond to edges incident to $v$, and the remaining one corresponds to the empty grid segment. Since the latter is a fixed source, it sends $1$ unit of flow to $s$; thus, the total inflow to $s$ from the former three sources must be $3$. Thus, every vertex has exactly one incoming edge (corresponding to flow value $2$) and exactly one outgoing edge (corresponding to flow value 0) in the constructed subgraph, forming a valid cycle cover. The two constructions are inverses of each other. $\blacksquare$
@@
Degree at least $3$. Incident to exactly one exclusive source and two unconstrained or biased sources along orthogonal directions. The remaining direction, if any, is occupied by a fixed source.
@@
Degree at least $2$. Incident to a specified pair of adjacent non-fixed sources: two unconstrained, two exclusive, or two biased (one inward and one outward). Any remaining directions are occupied by fixed sources.
@@
Let $G$ be a planar degree-$3$ bipartite mixed graph in which every vertex is incident to exactly one required edge. We can construct a flow network $N$ for $G$ using only T-Shaped, L-shaped, and blank sinks.
@@
Now, we construct such an embedding. Since each vertex has degree $3$, it forms a T-shape in the rectilinear embedding. Suppose that we place a vertex at an arbitrary lattice point. We align the incident segment corresponding to the required edge along one direction, and the segments corresponding to the other incident edges along the orthogonal directions. By choosing the orientation of each T-shape according to the coordinate parity, we can ensure that all required edges share one parity and all non-required edges share the other. Given this parity assignment, the actual embedding can be readily constructed by replacing each lattice point with a unit of $2\times 2$ lattice points as in Figure˜5.
@@
Finally, using this embedding, we construct the flow network $N$. Each vertex becomes a T-shaped sink, each intermediate point on an edge becomes an L-shaped sink of the corresponding type, and each empty lattice point becomes a blank sink. $\blacksquare$
@@ 4  ASP-Completeness of Constraint Graph Satisfiability
CGS for planar graphs is called Planar CGS. Typically, all vertex weights are restricted to $2$, and all edge weights are restricted to either $1$ or $2$. Each edge is colored according to its weight: weight-$1$ edges are red, and weight-$2$ edges are blue.
@@
Furthermore, edges are often allowed to have different weights at their endpoints; that is, an edge can have weight $2$ at one endpoint and weight $1$ at the other. Problems containing such edges are referred to as having arbitrary edge weights, while the standard case is referred to as having matching edge weights [MITHardness24GeneralizedSAT]. In this work, we focus on the latter.
@@
Proof.We construct a reduction from Restricted CCP. We first describe the construction without considering planarity. The core idea of the reduction is as follows: Each vertex of the CCP instance is colored either blue or red according to its bipartition; this coloring corresponds to the type of the associated vertex gadget. A vertex gadget has three terminals, corresponding to its three incident edges. Each terminal admits three possible states, denoted by $I$, $O$, and $U$, representing an incoming edge, an outgoing edge, and an unused edge in a cycle cover, respectively. Each vertex gadget admits exactly those assignments of terminal states in which each of $I$, $O$, and $U$ appears exactly once. An edge gadget connects a terminal of a blue vertex gadget to a terminal of a red vertex gadget. It represents a single wire that transmits the state between the two endpoints such that $I$ and $O$ are swapped, while $U$ remains unchanged. An undirected edge gadget admits all three possible states, whereas a directed edge gadget admits only two; it forbids the state corresponding to traversal against the original edge’s direction (i.e., $I$ at the tail terminal and $O$ at the head terminal). Under this construction, any valid assignment of states to the gadgets corresponds one-to-one to a cycle cover of the original instance.
We now describe the gadgets in detail. The vertex gadgets are illustrated in Figure˜7. Each terminal of a vertex gadget consists of three parallel edges labeled $I$, $O$, and $U$, representing the three possible states. As will be demonstrated, in a terminal of a blue (resp., red) vertex gadget, exactly one of these three edges must be oriented inward (resp., outward); the label of this edge determines the state of the terminal. With this convention, it is straightforward to verify that both types of vertex gadgets admit exactly those assignments of states described above.
The edge gadgets are illustrated in Figure˜8. Each gadget consists of three parallel paths connecting a blue vertex gadget to a red vertex gadget. Each end of these paths is labeled $I$, $O$, and $U$ according to the corresponding terminal of the vertex gadget. These paths connect the corresponding labels as follows: $I$ to $O$, $O$ to $I$, and $U$ to $U$. As we show below, exactly one of these three paths must be oriented from red to blue; the pair of labels of this path determines the state of the gadget. It can be verified that the undirected edge gadget (Figure˜8(a)) admits all three possible states, whereas the directed edge gadget (Figures˜8(b) and 8(c)) forbids the specified state.
@@ 5.1  Kakuro
    Given a grid, assign digits from $1$ to $9$ to each blank cell.
@@
Observe that for any $L^{\prime}_{\text{min}}\leq L_{\text{min}}\leq L_{\text{max}}\leq L^{\prime}_{\text{max}}$, every instance of $(N,L_{\text{min}},L_{\text{max}})$-Kakuro is also a valid instance of $(N,L^{\prime}_{\text{min}},L^{\prime}_{\text{max}})$-Kakuro. Hence, $(N,L_{\text{min}},L_{\text{max}})$-Kakuro is trivially reducible to $(N,L^{\prime}_{\text{min}},L^{\prime}_{\text{max}})$-Kakuro; in particular, it is parsimonious. It follows that, for fixed $N$, the hardness of $(N,L_{\text{min}},L_{\text{max}})$-Kakuro (or its counting version) implies the hardness of $(N,L^{\prime}_{\text{min}},L^{\prime}_{\text{max}})$-Kakuro for all $L^{\prime}_{\text{min}}\leq L_{\text{min}}$ and $L^{\prime}_{\text{max}}\geq L_{\text{max}}$. Similarly, the tractability of $(N,L^{\prime}_{\text{min}},L^{\prime}_{\text{max}})$-Kakuro (or its counting version) implies the tractability of $(N,L_{\text{min}},L_{\text{max}})$-Kakuro.
In contrast, for fixed $L_{\text{min}}$ and $L_{\text{max}}$, the complexity of $(N,L_{\text{min}},L_{\text{max}})$-Kakuro may vary with $N$. This is because increasing $N$ enlarges the set of available digits, thereby increasing the number of potential assignments to blank cells. Thus, although $(N,L_{\text{min}},L_{\text{max}})$-Kakuro is trivially reducible to $(N^{\prime},L_{\text{min}},L_{\text{max}})$-Kakuro for $N\leq N^{\prime}$, the reduction is not parsimonious in general.
@@
    $(N,2,2)$-Kakuro is in P for all $N$ [Seta02CrossSum].
@@
This result strengthens the previously known hardness results. Moreover, it is straightforward that #$(N,2,2)$-Kakuro is in FP for all $N$, since the input puzzle board must be a $2\times 2$ block. Consequently, this theorem allows us to completely classify the counting complexity of $(N,2,L_{\text{max}})$-Kakuro for all $N\geq 2$ and $L_{\text{max}}\geq 2$.
@@
Proof of Theorem˜14.We construct a reduction from Restricted CCP. Let $G=(V_{1}\cup V_{2},E,A)$ be a planar degree-3 bipartite mixed graph in which every vertex is incident to a directed edge. The key idea is to encode the state of each edge in a cycle cover of $G$ using digits assigned to the terminals of Kakuro gadgets. Specifically, an edge directed from a vertex in $V_{1}$ to a vertex in $V_{2}$ corresponds to digit $2$, an edge in the opposite direction corresponds to digit $3$, and an unused edge corresponds to digit $1$.
@@
Since every vertex in $G$ is incident to at least one directed edge, it suffices to construct two types of vertex gadgets per color: one incident to an incoming edge and one incident to an outgoing edge. Figure˜12 illustrates the former. The lower-left terminal represents an incoming edge and admits digits in $\{1,3\}$ in the red vertex gadget (resp., $\{1,2\}$ in the yellow vertex gadget). The latter is obtained by replacing all clues $3$ with $4$ (resp., $4$ with $3$). Note that a $90^{\circ}$ rotation exchanges terminal colors, and thereby reverses the direction of the edge.
@@
Vertex gadgets are connected using three types of edge gadgets: directed (red-to-yellow and yellow-to-red), and undirected. Figure˜13 illustrates the components of a directed edge gadget from red to yellow. Each component admits exactly two local solutions, forcing the terminal values to be either both $1$ or both $2$. We can extend the gadget, bend it, or invert its terminal parity. A directed edge gadget from yellow to red is obtained by swapping clue values $3$ and $4$. The components of an undirected edge gadget are illustrated in Figure˜14. Each component admits exactly three local solutions, corresponding to terminal values $1$, 22, or $3$. As with directed edges, these gadgets can be extended and bent. By appropriately configuring the vertex gadgets, we can match the parity of any terminal that admits an undirected edge; thus no parity inversion component is needed. These edge gadgets can, indeed, be placed so as not to interfere with one another or with vertex gadgets.
@@
Figure 13: Components of a directed edge gadget for Kakuro (from red to yellow) and an example of their local solutions. The two terminals of each component admit the same digit: either $1$ or $2$. The straight component can be extended by inserting additional blank cells with appropriate clues.
@@
Figure 14: Components of an undirected edge gadget for Kakuro and an example of their local solutions. The two terminals of each component admit the same digit: either 11, 22, or $3$. The straight component can be extended at the portion that has a similar structure to the straight component of a directed edge gadget.
The resulting Kakuro instance contains runs of length only $2$ and $3$, and any solution uses only digits in $\{1,2,3\}$ regardless of the availability of other digits. This completes the proof. $\blacksquare$

### 06 Mar 2026 v2

@@ 1  Introduction
Another framework is the Constraint Graph Satisfiability problem (CGS) introduced in [HearnDemaine09Puzzles, Section 5.1.3], which is a variant of the graph orientation problem defined on constraint graphs. In essence, a constraint graph is a graph with edge weights 11 or 22, where the vertices are of three types: and, or, and maj (majority). While planar CGS with only and and or vertices is known to be NP-complete [HearnDemaine09Puzzles, Theorem 5.4], extending this result to ASP-completeness is nontrivial. This is because the known reduction relies on a red-blue conversion gadget ([HearnDemaine09Puzzles, Figure 2.4]) that is not parsimonious, and constructing a parsimonious version is conjectured to be impossible [MITHardness24GeneralizedSAT]. To address this distinction, recent work [MITHardness24GeneralizedSAT] formalized two variants of CGS: arbitrary edge weights, where an edge may have different weights at its two endpoints, and the standard matching edge weights, where an edge must have the same weight at both endpoints; the former is essentially equivalent to a red-blue conversion. Regarding the restriction of vertex types, the complexity classification for CGS with arbitrary edge weights is now fully settled [MITHardness24GeneralizedSAT]; in particular, the case with and, or, and maj vertices is known to be ASP-complete even when the graph is planar. In contrast, the ASP-completeness of CGS with matching edge weights using and, or, and maj vertices has remained open.
@@ 1.1  Our Contributions
In this paper, we propose an ASP-completeness framework designed to exploit the rich structural properties of puzzles, and demonstrate its versatility by applying it to several puzzle types. Specifically, our framework is based on a variant of the Cycle Cover Problem (CCP), which asks whether a given graph can be covered by vertex-disjoint cycles. CCP on mixed graphs (i.e., graphs with both directed and undirected edges) was originally introduced to analyze the ASP-completeness of Kakuro (Cross Sum) puzzles, and was shown to be ASP-complete even when restricted to planar max-degree-33 graphs [Seta02CrossSum]. First, we introduce required edges—edges that must be included in the cycle cover—into CCP on mixed graphs, and define this problem as the Required-edge Cycle Cover Problem (RCCP). We focus on a subclass of instances on planar degree-33 bipartite graphs where each vertex is incident to exactly one required edge; we refer to this problem as Restricted RCCP. We prove that Restricted RCCP is ASP-complete. Moreover, by constructing a parsimonious gadget simulating a required edge, we obtain a corollary for CCP restricted to planar degree-33 bipartite mixed graphs where each vertex is incident to at least one directed edge. We denote this variant as Restricted CCP and conclude that it is also ASP-complete.
Second, we establish the ASP-completeness of several problems via reductions from Restricted RCCP/CCP. Crucial to these reductions is the exclusive structure of Restricted RCCP/CCP: since every vertex is of degree 33, fixing a cycle cover enforces a strict local constraint where each vertex must have exactly one incoming edge, one outgoing edge, and one unused edge. From the edge perspective, each edge can take one of three states: being traversed in either direction or remaining unused (with directed edges limited to two states). Since each consistent assignment of these states corresponds to a unique cycle cover, this structure allows us to simulate Restricted RCCP/CCP using simple vertex and edge gadgets. Furthermore, the bipartiteness of the graphs permits the use of asymmetric edge gadgets and facilitates the geometric embedding of these gadgets into the rectangular grids typical of puzzles.
@@ 2  Preliminaries
A mixed graph $G=(V,E,A)$ is a graph consisting of a set of vertices $V$, a set of undirected edges $E$, and a set of directed edges AA. The degree of a vertex $v$ is defined as the number of undirected edges incident to $v$ plus the number of directed edges incident to $v$ (i.e., having $v$ as a head or tail).
@@ 3.2  Flow Model Equivalent to Restricted RCCP
In order to construct the model, we assume that each source has an in-degree of 0, an out-degree of 22, and a supply of 22. Under this assumption, an unconstrained source admits any flow distributions in $\{(0,2),(1,1),(2,0)\}$. We additionally employ three variants that restrict the admissible flow distributions as follows:
@@
$\{(1,1),(2,0)\}$. Must send at least 11 unit along the first edge. We say such a source is oriented toward that edge.
@@
Formally, we transform an instance of Restricted RCCP into a flow network as follows. First, given a planar degree-33 mixed graph $G$ with required edges, we rectilinearly embed it into a grid (Figure˜4(a)). Let $H$ denote the resulting grid graph containing the embedding (Figure˜4(b)). We construct a network NN from $H$ based on the following rules (Figure˜4(c)):
@@
    Each segment (including empty ones) of $H$ is subdivided by inserting a source of supply 22; the two resulting edges are oriented toward the adjacent sinks.
@@
Now, we associate cycle covers of $G$ with flows in NN as follows. For each vertex $v$ and its corresponding sink, an edge incident to $v$ is incoming, outgoing, or unused in the cycle cover if and only if the corresponding flow value is 22, 0, or 11, respectively. Under this correspondence, we have the following lemma.
@@
lemmalemflowmodel Let $G$ be a planar degree-33 mixed graph in which every vertex is incident to exactly one required edge, and let NN be the flow network constructed as described above. Then there is a bijection between the cycle covers of $G$ and the flows in NN.
Proof.We first verify the forward direction. Fix a cycle cover of $G$, and consider its corresponding cycle family in $H$. Each lattice point in $H$ is either traversed by the cycle family or not.. In the former case, the sources corresponding to the incoming and outgoing edges of the cycle contribute 22 and 0 units of flow, respectively; each source on the remaining incident edges contributes 11 unit. In the latter case, all incident edges are unused; thus each source contributes 11 unit of flow. Therefore, the total inflow to the sink equals its degree in $H$ in any case.
Conversely, we show that a feasible flow in NN induces a valid cycle cover of $G$. Fix a feasible flow in NN. We construct a subgraph of $G$ and orient its edges based on the flow interpretation defined above. Flow conservation at intermediate lattice points ensures that each edge of $G$ has a well-defined orientation. Furthermore, due to the biased sources, this orientation is consistent with the direction of the directed edges in $G$.
The remaining task is to show that this subgraph forms a valid cycle cover. Consider a vertex $v$ in $G$ and its corresponding sink $s$ in NN. Assume first that $s$ has degree 33. Since $v$ has degree 33 and is incident to exactly one required edge, one of the three sources adjacent to $s$ is an exclusive source, which supplies either 0 or 22 units of flow. Since the flow from each source is at most 22, the only way that three nonnegative integers, one of which is in $\{0,2\}$, can sum to 33 is for the multiset of contributions to be $\{0,1,2\}$. The case where $s$ has degree 44 is similar. Indeed, exactly three of its adjacent sources correspond to edges incident to $v$, and the remaining one corresponds to the empty grid segment. Since the latter is a fixed source, it sends 11 unit of flow to $s$; thus, the total inflow to $s$ from the former three sources must be 33. Thus, every vertex has exactly one incoming edge (corresponding to flow value 22) and exactly one outgoing edge (corresponding to flow value 0) in the constructed subgraph, forming a valid cycle cover. The two constructions are inverses of each other. $\blacksquare$
@@
Degree at least 33. Incident to exactly one exclusive source and two unconstrained or biased sources along orthogonal directions. The remaining direction, if any, is occupied by a fixed source.
@@
Degree at least 22. Incident to a specified pair of adjacent non-fixed sources: two unconstrained, two exclusive, or two biased (one inward and one outward). Any remaining directions are occupied by fixed sources.
@@
Let $G$ be a planar degree-33 bipartite mixed graph in which every vertex is incident to exactly one required edge. We can construct a flow network NN for $G$ using only T-Shaped, L-shaped, and blank sinks.
@@
Now, we construct such an embedding. Since each vertex has degree 33, it forms a T-shape in the rectilinear embedding. Suppose that we place a vertex at an arbitrary lattice point. We align the incident segment corresponding to the required edge along one direction, and the segments corresponding to the other incident edges along the orthogonal directions. By choosing the orientation of each T-shape according to the coordinate parity, we can ensure that all required edges share one parity and all non-required edges share the other. Given this parity assignment, the actual embedding can be readily constructed by replacing each lattice point with a unit of $2\times 2$ lattice points as in Figure˜5.
@@
Finally, using this embedding, we construct the flow network NN. Each vertex becomes a T-shaped sink, each intermediate point on an edge becomes an L-shaped sink of the corresponding type, and each empty lattice point becomes a blank sink. $\blacksquare$
@@ 4  ASP-Completeness of Constraint Graph Satisfiability
CGS for planar graphs is called Planar CGS. Typically, all vertex weights are restricted to 22, and all edge weights are restricted to either 11 or 22. Each edge is colored according to its weight: weight-11 edges are red, and weight-22 edges are blue.
@@
Furthermore, edges are often allowed to have different weights at their endpoints; that is, an edge can have weight 22 at one endpoint and weight 11 at the other. Problems containing such edges are referred to as having arbitrary edge weights, while the standard case is referred to as having matching edge weights [MITHardness24GeneralizedSAT]. In this work, we focus on the latter.
@@
Proof.We construct a reduction from Restricted CCP. We first describe the construction without considering planarity. The core idea of the reduction is as follows: Each vertex of the CCP instance is colored either blue or red according to its bipartition; this coloring corresponds to the type of the associated vertex gadget. A vertex gadget has three terminals, corresponding to its three incident edges. Each terminal admits three possible states, denoted by II, $O$, and $U$, representing an incoming edge, an outgoing edge, and an unused edge in a cycle cover, respectively. Each vertex gadget admits exactly those assignments of terminal states in which each of II, $O$, and $U$ appears exactly once. An edge gadget connects a terminal of a blue vertex gadget to a terminal of a red vertex gadget. It represents a single wire that transmits the state between the two endpoints such that II and $O$ are swapped, while $U$ remains unchanged. An undirected edge gadget admits all three possible states, whereas a directed edge gadget admits only two; it forbids the state corresponding to traversal against the original edge’s direction (i.e., II at the tail terminal and $O$ at the head terminal). Under this construction, any valid assignment of states to the gadgets corresponds one-to-one to a cycle cover of the original instance.
We now describe the gadgets in detail. The vertex gadgets are illustrated in Figure˜7. Each terminal of a vertex gadget consists of three parallel edges labeled II, $O$, and $U$, representing the three possible states. As will be demonstrated, in a terminal of a blue (resp., red) vertex gadget, exactly one of these three edges must be oriented inward (resp., outward); the label of this edge determines the state of the terminal. With this convention, it is straightforward to verify that both types of vertex gadgets admit exactly those assignments of states described above.
The edge gadgets are illustrated in Figure˜8. Each gadget consists of three parallel paths connecting a blue vertex gadget to a red vertex gadget. Each end of these paths is labeled II, $O$, and $U$ according to the corresponding terminal of the vertex gadget. These paths connect the corresponding labels as follows: II to $O$, $O$ to II, and $U$ to $U$. As we show below, exactly one of these three paths must be oriented from red to blue; the pair of labels of this path determines the state of the gadget. It can be verified that the undirected edge gadget (Figure˜8(a)) admits all three possible states, whereas the directed edge gadget (Figures˜8(b) and 8(c)) forbids the specified state.
@@ 5.1  Kakuro
    Given a grid, assign digits from 11 to 99 to each blank cell.
@@
Observe that for any $L^{\prime}_{\text{min}}\leq L_{\text{min}}\leq L_{\text{max}}\leq L^{\prime}_{\text{max}}$, every instance of $(N,L_{\text{min}},L_{\text{max}})$-Kakuro is also a valid instance of $(N,L^{\prime}_{\text{min}},L^{\prime}_{\text{max}})$-Kakuro. Hence, $(N,L_{\text{min}},L_{\text{max}})$-Kakuro is trivially reducible to $(N,L^{\prime}_{\text{min}},L^{\prime}_{\text{max}})$-Kakuro; in particular, it is parsimonious. It follows that, for fixed NN, the hardness of $(N,L_{\text{min}},L_{\text{max}})$-Kakuro (or its counting version) implies the hardness of $(N,L^{\prime}_{\text{min}},L^{\prime}_{\text{max}})$-Kakuro for all $L^{\prime}_{\text{min}}\leq L_{\text{min}}$ and $L^{\prime}_{\text{max}}\geq L_{\text{max}}$. Similarly, the tractability of $(N,L^{\prime}_{\text{min}},L^{\prime}_{\text{max}})$-Kakuro (or its counting version) implies the tractability of $(N,L_{\text{min}},L_{\text{max}})$-Kakuro.
In contrast, for fixed $L_{\text{min}}$ and $L_{\text{max}}$, the complexity of $(N,L_{\text{min}},L_{\text{max}})$-Kakuro may vary with NN. This is because increasing NN enlarges the set of available digits, thereby increasing the number of potential assignments to blank cells. Thus, although $(N,L_{\text{min}},L_{\text{max}})$-Kakuro is trivially reducible to $(N^{\prime},L_{\text{min}},L_{\text{max}})$-Kakuro for $N\leq N^{\prime}$, the reduction is not parsimonious in general.
@@
    $(N,2,2)$-Kakuro is in P for all NN [Seta02CrossSum].
@@
This result strengthens the previously known hardness results. Moreover, it is straightforward that #$(N,2,2)$-Kakuro is in FP for all NN, since the input puzzle board must be a $2\times 2$ block. Consequently, this theorem allows us to completely classify the counting complexity of $(N,2,L_{\text{max}})$-Kakuro for all $N\geq 2$ and $L_{\text{max}}\geq 2$.
@@
Proof of Theorem˜14.We construct a reduction from Restricted CCP. Let $G=(V_{1}\cup V_{2},E,A)$ be a planar degree-3 bipartite mixed graph in which every vertex is incident to a directed edge. The key idea is to encode the state of each edge in a cycle cover of $G$ using digits assigned to the terminals of Kakuro gadgets. Specifically, an edge directed from a vertex in $V_{1}$ to a vertex in $V_{2}$ corresponds to digit 22, an edge in the opposite direction corresponds to digit 33, and an unused edge corresponds to digit 11.
@@
Since every vertex in $G$ is incident to at least one directed edge, it suffices to construct two types of vertex gadgets per color: one incident to an incoming edge and one incident to an outgoing edge. Figure˜12 illustrates the former. The lower-left terminal represents an incoming edge and admits digits in $\{1,3\}$ in the red vertex gadget (resp., $\{1,2\}$ in the yellow vertex gadget). The latter is obtained by replacing all clues 33 with 44 (resp., 44 with 33). Note that a $90^{\circ}$ rotation exchanges terminal colors, and thereby reverses the direction of the edge.
@@
Vertex gadgets are connected using three types of edge gadgets: directed (red-to-yellow and yellow-to-red), and undirected. Figure˜13 illustrates the components of a directed edge gadget from red to yellow. Each component admits exactly two local solutions, forcing the terminal values to be either both 11 or both 22. We can extend the gadget, bend it, or invert its terminal parity. A directed edge gadget from yellow to red is obtained by swapping clue values 33 and 44. The components of an undirected edge gadget are illustrated in Figure˜14. Each component admits exactly three local solutions, corresponding to terminal values 11, 22, or 33. As with directed edges, these gadgets can be extended and bent. By appropriately configuring the vertex gadgets, we can match the parity of any terminal that admits an undirected edge; thus no parity inversion component is needed. These edge gadgets can, indeed, be placed so as not to interfere with one another or with vertex gadgets.
@@
Figure 13: Components of a directed edge gadget for Kakuro (from red to yellow) and an example of their local solutions. The two terminals of each component admit the same digit: either 11 or 22. The straight component can be extended by inserting additional blank cells with appropriate clues.
@@
Figure 14: Components of an undirected edge gadget for Kakuro and an example of their local solutions. The two terminals of each component admit the same digit: either 11, 22, or 33. The straight component can be extended at the portion that has a similar structure to the straight component of a directed edge gadget.
The resulting Kakuro instance contains runs of length only 22 and 33, and any solution uses only digits in $\{1,2,3\}$ regardless of the availability of other digits. This completes the proof. $\blacksquare$